OpenStudy (anonymous):

Twice the sum of a number and 3 times a second number is 4. The difference of ten times the second number and five times the first is 90. Find the numbers!!!

OpenStudy (agent0smith):

Try writing it the way we solved the last one.

OpenStudy (agent0smith):

ie follow the steps from last time.

OpenStudy (anonymous):

I did, idk how to do it

OpenStudy (anonymous):

Ley x = first number Let y = second number 2(x + 3y) = 4 10y - 5x = 90 2x + 6y = 4 -5x + 10y = 90 10x + 30y = 20 -10x + 20y = 180 50y = 200 y = 4 If y = 4, then substitute y = 4 in the first equation, 2(x + 3y) = 4; 2(x + 12) = 4 So, x + 12 = 2, and x = -10 The first number is 10 and the second number is 4.

OpenStudy (anonymous):

I messed up on the parenthesis, thanks:)